1 + 1 over x = 72 over x^2
Answer by fractalier(6550)

1 + (1/x) = 72/x^2
multiply all terms by x^2 and get
x^2 + x = 72
x^2 + x - 72 = 0
(x+9)(x-8) = 0
x = -9 and x = 8